Twice the product of two numbers is equal to the sum of their squares. What is the difference between the two numbers? And what are the numbers?
Answer by ankor@dixie-net.com(22740)

let the two numbers be a & b
Write an equation for exactly what it says
:
Twice the product of two numbers is equal to the sum of their squares.
2(ab) = a^2 + b^2
Arrange as a quadratic equation
a^2 - 2ab + b^2 = 0
Factors to
(a - b) (a - b) = 0
:
What is the difference between the two numbers?
the difference is 0
a = b
:
And what are the numbers
a & b can equal any number, 2ab is the same as a^2+b^2, when they are equal
: